Ann is now seven times as old as Joan. In six years, Ann will be five times as old as Joan. Find the present age of each one.
Answer by stanbon(75887)

Ann is now seven times as old as Joan. In six years, Ann will be five times as old as Joan. Find the present age of each one.
-----------------
Equations:
A = 7J
(A+6)=5(J+6)
---
Substitute for A and solve for "J":
(7J+6) = 5(J+6)
2J = 24
J = 12 (Joan's age now)
Since A = 7J, A = 84 (Ann's age now)
========================================
Cheers,
Stan H.